The total product of the 10 employees of ABC is 10,000 bags . the firm is considering employing a Chinese worker whose marginal product is 900 bags .each bags is sold at K10 and all the workers are paid a flat wage of K8,900



A. What is the average product before and after this workers is employed?



B. Should ABC employ the new Chinese worker?



C. What is the minimum marginal product that ABC can accept for any additional worker under the prevailing price and wage rate?

1
Expert's answer
2021-11-19T11:13:06-0500

A.

"AP=\frac{TP}{L}"

Before:

"AP=\frac{10,000}{10}=1,000"

After:

"MP=900"

"∆Q=900\times1=900"

"AP=\frac{10,900}{11}=991"


B.

ABC should not employ the new Chinese worker because average product of ABC declines implying that the rate of production return per unit produced after addition of the Chinese worker is diminishing.